x^2 -6x +14

x^2 -6x +14

Grade:12

2 Answers

AskIITians Expert Hari Shankar IITD
14 years ago

x2 - 6x + 14 = 0

Discriminant D = b2-4ac

= 62 - 4(1)(14) = 36 - 56 = -20

Since the Discriminant D is negative, this equation has no real solutions. It has only imaginary solutions.

square root(D) = sqrt(-20) = 2i sqrt(5)

x= (-b + sqrt(D))/2a , (-b - sqrt(D))/2a

or, x = (6 + 2i sqrt(5))/2 , (6- 2i sqrt(5))/2

or, x = 3 + i sqrt(5), 3 - i sqrt(5) are the imaginary solutions.

 

(If you have not studied complex numbers, leave out the part after the 4th line)

anurag singhal
14 years ago

this equation does not have real roots

D<0

it will be 6+/_iroot20